OpenStudy (anonymous):

Write the simplest polynomial with the zeros 2-i, Square root 5, and -2

OpenStudy (amoodarya):

f(x)=(x^2-5)(x-(-2))((x-2+i)((x-2-i) if coefficient are integers

OpenStudy (mathstudent55):

I assume you are looking for a polynomial with real coefficients. If that's the case, then a polynomial with real coefficients can only have complex roots if they come in complex conjugate pairs. That means, by having complex root 2 - i, you also have root 2 + i. In addition, you are told you also have roots \(\sqrt5\) and -2. That means you have the following 4 factors: \(y = (x - (2 - i))(x - (2 + i))(x - \sqrt{5})(x - (-2)) \) \( y = (x - 2 + i)(x - 2 - i)(x - \sqrt{5})(x + 2) \)